# Grover Speedup from Many Forms of the Zeno Effect

### v4

@@ 3.1 Applying our Strategy to Quantum Walk and Adiabatic settings
It is further clear that except for very special values of $\gamma$ and $t$, the final overlap with $|m\rangle$ is non-zero, however, making sure this value does not vanish as $N$ is scaled is key to demonstrate a speedup. A quantum walk at exactly the avoided crossing or within a fixed finite multiple of $g_{\mathrm{min}}$ of the avoided crossing will always attain an optimal quantum speedup, but this requires exponentially increasing precision.
@@
The structure of the argument makes it clear that within this rescaled picture, a wide variety of schedules will yield an optimal speedup. In fact, such a schedule need only have $f(\tau)$ finite for a finite fraction of $\tau$ values, and to avoid special cases where exact destructive interference could occur. This is a more general restatement of the result of [37]. The crucial ingredient for optimal speedup is for the schedule to scale in the correct way around the avoided crossing as the problem grows, other details are not particularly important.
@@ 4.2 Zeno Effect Induced by Decoherence
The simplest way (which as we show does not work) one could consider taking this model into a continuum limit is to consider applying decoherence of this form for an infinite number of vanishingly small periods of time, to do this we first expand the action of the decoherence over a short time period
@@
$\kappa=\frac{\Delta p_{\pm}}{\Delta t}\approx\frac{q}{t_{\mathrm{scale}}}\left|\frac{\langle e|-\langle g|}{\sqrt{2}}\exp\left(\frac{-it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{m}\tilde{Z}\right)\frac{|e\rangle+|g\rangle}{\sqrt{2}}\right|^{2}$
@@
Here we have effectively defined a Zeno time for our decoherence interaction [53], in our case the Zeno time for the deocoherence inducing interaction is
@@
As before this argument does not strictly apply to cases where a Zeno effect is manifested, since these systems will remain in an approximately pure state throughout, we examine these cases numerically to show that indeed a Zeno effect can lead to finding the marked state with nearly $100\%$ probability. Figure 7 shows that a continuous Zeno effect does indeed lead to success probability which approaches $100\%$. We also show an alternative model for decoherence in appendix F.
@@ 5.1 Zeno Effect from Partially Destructive Measurements
$\rho\rightarrow|g\rangle\!\langle g|\rho|g\rangle\!\langle g|+|d\rangle\!\langle e|\rho|e\rangle\!\langle d|=|g\rangle\!\langle g|\left\langle g\middle|\rho\middle|g\right\rangle+|d\rangle\!\langle d|\left\langle e\middle|\rho\middle|e\right\rangle.$ (29)
@@
The algorithm given in [47] will also work for this kind of measurement. A partially destructive measurement of this type right at the avoided crossing would result in the $|d\rangle$ state $50\%$ of the time, but would be found in the $|g\rangle$ state the other $50\%$, and of these cases a subsequent measurement would find $|m\rangle$ $50\%$ of the time, for an overall success probability of $25\%$ each time the algorithm is run, since this is still $O(1)$ then such an algorithm would still yield an optimal speedup. Similar arguments apply for measurements anywhere close to the avoided crossing. However, since amplitude in the excited state is “lost” this case does not have the elegant argument about purity which could be made in the projective case. In particular, a pair of sequential measurements where $|\gamma|$ takes the same large value, but has the opposite sign will result in the system being found completely in the $|d\rangle$ state since $|e\rangle$ with $\gamma$ would be $|g\rangle$ with $-\gamma$ and vice versa.
Zeno effects are however still possible, and indeed as the number of measurements are increased can approach $100\%$ success, as can be seen in figure 8. As with the case of projective measurement, there is no way to directly take a continuum limit of the partially destructive measurements, we instead need to take a continuous quantum channel which behaves like a partially destructive measurement, which as we discuss in the next section is dissipation out of the $|e\rangle$ state.
@@ 5.2 Zeno Effect from Dissipation
$=\left[|g\rangle\!\langle g|\exp\left(it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}p\right)+|e\rangle\!\langle e|\exp\left(-it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}p\right)\right]\otimes\mathbb{1}_{2}.$ (31)
@@
Let us now consider an additional single qubit which starts in state $|0\rangle$ can be evolved with the following Hamiltonian which couples it to the continuous variable
@@
We have now shown that discrete steps which perform dissipation from the $|e\rangle$ to a useless state $|d\rangle$ are physically realistic, numerical plots further show that these can give finite probability to be found in $|m\rangle$ and therefore yield an optimal speedup.
@@
Naively, from the equations we have written previously, it would appear that this dissipation model could yield a speedup which is even faster than the $\sqrt{N}$ Grover speedup. Namely, if we were able to take $t_{\mathrm{rot}}\propto N^{\frac{1}{4}}$ and $t_{\mathrm{scale}}\propto N^{\frac{1}{4}}$, we could obtain the same dynamics, but in a time which scales only as $N^{\frac{1}{4}}$. The reason that such an implementation is unphysical (it has to be, since it gives a scaling which is not possible in real quantum systems) is somewhat subtle. If we scaled in this way, $x_{mag}$ scales as $N^{-\frac{1}{4}}$. In such a scenario, the width of the initial state of the continuous variable which we have assumed are “sufficiently narrow” to be approximated as classical particles, would have to also scale down at least as $N^{-\frac{1}{4}}$, such scaling is physically unrealistic, and without it, the final wavepackets would increasingly overlap as $N$ scales, for a large $N$ it would not be possible to reliably distinguish $|e\rangle$ from $|g\rangle$.
@@
We now consider what happens in the continuum limit. As with the decoherence case, we can see that as $t_{\mathrm{scale}}$ is decreased the probability of measuring in $|d\rangle$ after applying $D(\tau,t_{\mathrm{scale}}/q,t_{\mathrm{rot}})$ will scale as $1-\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)\propto t^{2}_{\mathrm{scale}}/q^{2}$. Therefore, if we divide the evolution into $q$ segments with $q\rightarrow\infty$, no dynamics will happen unless we also scale $t_{\mathrm{scale}}\propto q$. The reason for this is actually another Zeno effect, we are measuring the rotated qubit with increasing frequency, eventually stopping its dynamics. This behaviour is illustrated numerically in figure 5 as hollow squares.
@@
In the limit described above, infinite $m$ and large but finite $q$, each stage measurement effectively describes a fraction of the probability to be in state $|e\rangle$ being moved to $|d\rangle$ (represented by aborting the computation in the event of a $|1\rangle$ measurement). For a single measurement, the probability of this happening is $\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)$ according to equation 34. For large $q$ this can be converted into a decay rate by dividing by time
@@ 7 Discussion
$U(t)\rho U^{\dagger}(t)$ $\frac{1}{2}U(t)\rho U^{\dagger}(t)+$ $\cos^{2}(\frac{1}{2}g(\tau)t)\rho+$
$\frac{1}{2}U(-t)\rho U^{\dagger}(-t)$ $\sin^{2}(\frac{1}{2}g(\tau)t)\Pi_{g}\rho\Pi_{g}$
@@
$\frac{\partial}{\partial t}\rho=$ $\frac{\partial}{\partial t}\rho=$ $\frac{\partial}{\partial t}\rho=$
$i\frac{g(\tau)}{2}\left[\rho\tilde{Z}-\tilde{Z}\rho\right]$ $\kappa(\tau)\left[\tilde{Z}\rho\tilde{Z}-\rho\right]$ $\kappa(\tau)\left[\Pi_{de}\rho\Pi_{ed}-\frac{1}{2}(\Pi_{e}\rho+\rho\Pi_{e})\right]$
@@
While the results we have shown are interesting from a theoretical perspective, a question we have not fully addressed thus far is what practical relevance the models we have discussed here have. In particular, while theoretically useful, all the Hamiltonians we have discussed here would be more complicated to implement than directly implementing an unstructured search Hamiltonian (for example using the methods of [61]). A natural question is whether there is actually a physical situation where it may be easier to implement either a dephasing or destruction model. We discuss at least one technology which answers this question in the affirmative for the dissipation model.
@@
In this setting, the Zeno effect is modulated by the off diagonal density matrix element which captures the coherence between a freely propagating mode (corresponding to $\hat{a}_{f}$) and the captured cavity mode (corresponding to $\hat{a}_{\mathrm{cav}}$). In fact, only the imaginary component of this term contributes to transfer of a freely propagating photon into the cavity, this leads to (assuming $m$ photons are present in the control mode and $n$ in the freely propagating mode),
@@
which is the same underlying differential equation as a damped harmonic oscillator. The term proportional to the loss rate $\gamma$ acts as an effective damping term, When $4\,\sqrt{n\,m}\,G>\gamma$, this equation is effectively under damped and $\mathrm{Im}\left[\rho_{cf}\right]$ will oscillate between positive and negative values, as it decays, characteristic in a phase rotation manifestation of the Zeno effect. On the other hand if $4\,\sqrt{n\,m}\,G\leq\gamma$ the system is over (or critically) damped and $\mathrm{Im}\left[\rho_{cf}\right]$, then this element will decay but not change sign, as would occur in a destruction based manifestation of the Zeno effect. These regimes roughly map to the coherent (CQZ) and incoherent (IQZ) quantum Zeno blockade regimes as discussed in previous literature [13, 62, 65] This regime is also the one which is more likely to be realised with small $m$ and $n$, in particular in the $m=n=1$ case of a single-photon Zeno blockade. This few or single photon regime is the one which is most relevant to entropy computing [12].
@@ Appendix A Review of the model from [47]
However, we must recall a key property of measurement is continuous systems is that the position cannot be measured with infinite precision (both because $x$ eigenstates cannot be prepared perfectly, and because of limitations of the apparatus, we consider the former in the main text for convenience). This precision is clearly independent of any properties of $\hat{A}$, or the runtime of the experiment. If we assume a fixed precision at which $x$ measurements can be reliably performed $\Delta x$, and our aim is to differentiate states with eigenvalues $a_{0}<a_{1}$, then the required time for the measurement will be
@@
in other words, it will depend inversely on the difference in expectation values of $\hat{A}$, which is the energy gap in the case where $\hat{A}$ is a Hamiltonian.

### v5

@@ 3.1 Applying our Strategy to Quantum Walk and Adiabatic settings
It is further clear that except for very special values of $\gamma$ and $t$, the final overlap with $|m\rangle$ is non-zero; however, making sure this value does not vanish as $N$ is scaled is key to demonstrate a speedup. A quantum walk at exactly the avoided crossing or within a fixed finite multiple of $g_{\mathrm{min}}$ of the avoided crossing will always attain an optimal quantum speedup, but this requires exponentially increasing precision.
@@
The structure of the argument makes it clear that within this rescaled picture, a wide variety of schedules will yield an optimal speedup. In fact, such a schedule need only have $f(\tau)$ finite for a finite fraction of $\tau$ values, and to avoid special cases where exact destructive interference could occur. This is a more general restatement of the result of [42]. The crucial ingredient for optimal speedup is for the schedule to scale in the correct way around the avoided crossing as the problem grows. Other details are not particularly important.
@@ 4.2 Zeno Effect Induced by Decoherence
The simplest way (which as we show does not work) one could consider taking this model into a continuum limit is to consider applying decoherence of this form for an infinite number of vanishingly small periods of time. To do this we first expand the action of the decoherence over a short time period
@@
$\kappa=\frac{\Delta p_{\pm}}{\Delta t}\approx\frac{q}{t_{\mathrm{scale}}}\left|\frac{\langle e|-\langle g|}{\sqrt{2}}\exp\left(\frac{-it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}}{q}\tilde{Z}\right)\frac{|e\rangle+|g\rangle}{\sqrt{2}}\right|^{2}$
@@
Here we have effectively defined a Zeno time for our decoherence interaction [58], in our case the Zeno time for the decoherence inducing interaction is
@@
As before this argument does not strictly apply to cases where a Zeno effect is manifested, since these systems will remain in an approximately pure state throughout. We examine these cases numerically to show that indeed a Zeno effect can lead to finding the marked state with nearly $100\%$ probability. Figure 7 shows that a continuous Zeno effect does indeed lead to success probability which approaches $100\%$. We also show an alternative model for decoherence in appendix F.
@@ 5.1 Zeno Effect from Partially Destructive Measurements
$\rho\rightarrow|g\rangle\!\langle g|\rho|g\rangle\!\langle g|+|d\rangle\!\langle e|\rho|e\rangle\!\langle d|+|d\rangle\!\langle d|\rho|d\rangle\!\langle d|=|g\rangle\!\langle g|\left\langle g\middle|\rho\middle|g\right\rangle+|d\rangle\!\langle d|\left(\left\langle e\middle|\rho\middle|e\right\rangle+\left\langle d\middle|\rho\middle|d\right\rangle\right).$ (29)
@@
The algorithm given in [52] will also work for this kind of measurement. A partially destructive measurement of this type right at the avoided crossing would result in the $|d\rangle$ state $50\%$ of the time, but would be found in the $|g\rangle$ state the other $50\%$, and of these cases a subsequent measurement would find $|m\rangle$ $50\%$ of the time, for an overall success probability of $25\%$ each time the algorithm is run. Since this is still $O(1)$ such an algorithm would still yield an optimal speedup. Similar arguments apply for measurements anywhere close to the avoided crossing. However, since amplitude in the excited state is “lost” this case does not have the elegant argument about purity which could be made in the projective case. In particular, a pair of sequential measurements where $|\gamma|$ takes the same large value, but has the opposite sign will result in the system being found completely in the $|d\rangle$ state since $|e\rangle$ with $\gamma$ would be $|g\rangle$ with $-\gamma$ and vice versa.
Zeno effects are however still possible, and indeed as the number of measurements are increased can approach $100\%$ success, as can be seen in figure 8. As with the case of projective measurement, there is no way to directly take a continuum limit of the partially destructive measurements. We instead need to take a continuous quantum channel which behaves like a partially destructive measurement, which as we discuss in the next section is dissipation out of the $|e\rangle$ state.
@@ 5.2 Zeno Effect from Dissipation
$=\left[|g\rangle\!\langle g|\exp\left(it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}p\right)+|e\rangle\!\langle e|\exp\left(-it_{\mathrm{scale}}\sqrt{f^{2}(\tau)+1}p\right)\right]\otimes\mathbb{1}_{2},$ (31)
where the purpose of the additional qubit subspace will become clear later.
@@
Let us now consider the additional single qubit which starts in state $|0\rangle$ can be evolved with the following Hamiltonian which couples it to the continuous variable
@@
We have now shown that discrete steps which perform dissipation from the $|e\rangle$ to a useless state $|d\rangle$ are physically realistic. Numerical plots further show that these can give finite probability to be found in $|m\rangle$ and therefore yield an optimal speedup.
@@
Naively, from the equations we have written previously, it would appear that this dissipation model could yield a speedup which is even faster than the $\sqrt{N}$ Grover speedup. Namely, if we were able to take $t_{\mathrm{rot}}\propto N^{\frac{1}{4}}$ and $t_{\mathrm{scale}}\propto N^{\frac{1}{4}}$, we could obtain the same dynamics, but in a time which scales only as $N^{\frac{1}{4}}$. The reason that such an implementation is unphysical (it has to be, since it gives a scaling which is not possible in real quantum systems) is somewhat subtle. If we scaled in this way, $x_{mag}$ scales as $N^{-\frac{1}{4}}$. In such a scenario, the width of the initial state of the continuous variable which we have assumed are “sufficiently narrow” to be approximated as classical particles, would have to also scale down at least as $N^{-\frac{1}{4}}$. Such scaling is physically unrealistic, and without it, the final wavepackets would increasingly overlap as $N$ scales. For a large $N$ it would not be possible to reliably distinguish $|e\rangle$ from $|g\rangle$.
@@
We now consider what happens in the continuum limit. As with the decoherence case, we can see that as $t_{\mathrm{scale}}$ is decreased the probability of measuring in $|d\rangle$ after applying $D(\tau,t_{\mathrm{scale}}/q,t_{\mathrm{rot}})$ will scale as $1-\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)\propto t^{2}_{\mathrm{scale}}/q^{2}$. Therefore, if we divide the evolution into $q$ segments with $q\rightarrow\infty$, no dynamics will happen unless we also scale $t_{\mathrm{scale}}\propto q$. The reason for this is actually another Zeno effect. We are measuring the rotated qubit with increasing frequency, eventually stopping its dynamics. This behaviour is illustrated numerically in figure 5 as hollow squares.
@@
In the limit described above, infinite $m$ and large but finite $q$, each measurement effectively describes a fraction of the probability to be in state $|e\rangle$ being moved to $|d\rangle$ (represented by aborting the computation in the event of a $|1\rangle$ measurement). For a single measurement, the probability of this happening is $\sin^{2}\left(2t_{\mathrm{rot}}x_{\mathrm{mag}}\right)$ according to equation 34. For large $q$ this can be converted into a decay rate by dividing by time
@@ 7 Discussion
[…]
@@
[…]
@@
While the results we have shown are interesting from a theoretical perspective, a question we have not fully addressed thus far is what practical relevance the models we have discussed here have. In particular, while theoretically useful, all the Hamiltonians we have discussed here would be more complicated to implement than directly implementing an unstructured search Hamiltonian (for example using the methods of [66]). A natural question is whether there is actually a physical situation where it may be easier to implement either a dephasing or destruction model. We discuss at least one technology which answers this question in the affirmative for the destruction model. Note that we will not discuss how to actually encode the problem to this paradigm, as that is beyond the scope of this current work.
@@
In this setting, the Zeno effect is modulated by the off-diagonal density matrix element which captures the coherence between a freely propagating mode (corresponding to $\hat{a}_{f}$) and the captured cavity mode (corresponding to $\hat{a}_{\mathrm{cav}}$). In fact, only the imaginary component of this term contributes to transfer of a freely propagating photon into the cavity, this leads to (assuming $m$ photons are present in the control mode and $n$ in the freely propagating mode),
@@
which is the same underlying differential equation as a damped harmonic oscillator. The term proportional to the loss rate $\gamma$ acts as an effective damping term, When $4\,\sqrt{n\,m}\,G>\gamma$, this equation is effectively under damped and $\mathrm{Im}\left[\rho_{cf}\right]$ will oscillate between positive and negative values, as it decays, characteristic in a phase rotation manifestation of the Zeno effect. On the other hand if $4\,\sqrt{n\,m}\,G\leq\gamma$ the system is over (or critically) damped then $\mathrm{Im}\left[\rho_{cf}\right]$ will decay but not change sign, as would occur in a destruction based manifestation of the Zeno effect. These regimes roughly map to the coherent (CQZ) and incoherent (IQZ) quantum Zeno blockade regimes as discussed in previous literature [13, 67, 70] This regime is also the one which is more likely to be realised with small $m$ and $n$, in particular in the $m=n=1$ case of a single-photon Zeno blockade. This few or single photon regime is the one which is most relevant to entropy computing [12].
@@ Appendix A Review of the model from [52]
However, we must recall a key property of measurement of continuous systems is that the position cannot be measured with infinite precision (both because $x$ eigenstates cannot be prepared perfectly, and because of limitations of the apparatus, we consider the former in the main text for convenience). This precision is clearly independent of any properties of $\hat{A}$, or the runtime of the experiment. If we assume a fixed precision at which $x$ measurements can be reliably performed $\Delta x$, and our aim is to differentiate states with eigenvalues $a_{0}<a_{1}$, then the required time for the measurement will be
@@
in other words, it will depend inversely on the difference in expectation values of $\hat{A}$, which is the energy gap in the case where $\hat{A}$ is a Hamiltonian [^5] Note that as a Hamiltonian $\hat{A}$ would have a natural energy scale so a trivial rescaling of $\hat{A}\rightarrow c\hat{A}$ would not be allowed..